# A quantitative assessment of Geant4 for predicting the yield and distribution of positron-emitting fragments in ion beam therapy

### v1

@@ 3.4 Overall recommendation
In the entrance region, BIC was clearly the best-performing model, with the best choice of Geant4 version depending on the particular metric and fragmentation product of interest. NMSE results generally favoured 10-10.4/BIC (especially 10.2/BIC), except for ${}^{10}$ yield, which was better predicted by 10.3+/INCL. Pearson CCC performance did not strongly favour any particular version/model combination, with at most 1/3 of test cases achieving the target CCC threshold of 0.95 for any version/model.
In the build-up and Bragg peak region and tail region, the results are more clear-cut. Results for the NMSE metric conclusively show that version 10.2/BIC is clearly the best choice for total positron yield, ${}^{11}$ production and ${}^{15}$ yield, while 10.5-11.1/INCL performed the best for ${}^{10}$. Pearson CCC results are more mixed, but again, 10.2/BIC gives the best results for total positron annihilation yield, with most versions of Geant4 with BIC performing well. 10.6/QMD performed the best for ${}^{11}$, 10/INCL and 10.1/INCL performed the best for ${}^{10}$, and there was no clear winner for ${}^{15}$.
Using the depth-of-maximum-yield metric, the smallest mean differences were obtained with 10.4/QMD and 10.6/QMD. These versions/models also achieved the equal-smallest maximum difference (-3 mm and +3 mm, respectively). Across all versions of Geant4, QMD demonstrated the best overall accuracy (lowest average mean difference in peak depth) and highest precision (lowest average standard deviation). INCL was the worst-performing model across all versions, with much larger maximum differences, and a consistent underestimation of depth of maximum yield across Geant4 versions, with the exception of version 11.1 (which, despite a mean difference of 0, exhibited a large standard deviation and maximum value). Standard deviations obtained using INCL were generally around double those of QMD and BIC. BIC also showed a consistent underestimation in depth of maximum yield, although the maximum differences were much smaller than for INCL. For context, the difference between the depth of the positron annihilation peak and the Bragg peak with monoenergetic ion beams is of the order of -5.6$\pm$0.8 mm for ${}^{12}$ and -6.6$\pm$0.8 mm for ${}^{16}$ [32, 33, 6].
@@
Finally, in the tail region, Geant4 10.2 with BIC and QMD again provided the best prediction of total positron and ${}^{11}$ yield in terms of NMSE, while 10.6/INCL performed the best for ${}^{10}$ and ${}^{15}$. All version/model combinations performed well for total positron annihilation, ${}^{11}$ and ${}^{15}$ yield according to the Pearson CCC metric, while no version/model performed especially well for ${}^{10}$.
@@
In the evaluation of individual positron-emitting fragment yield profiles, predictions of ${}^{10}$ distribution were generally the least accurate in terms of both the NMSE and Pearson CCC. Interestingly, the INCL model often performed the best for prediction of ${}^{10}$ fragment yield, although it rarely performed the best for total positron annihilation and ${}^{11}$ or ${}^{15}$. Therefore, INCL should be considered for studies focusing on ${}^{10}$ fragmentation, with the caveat that range estimation will be less accurate with this model.
@@
Finally, it is worth noting that current evaluations of fragmentation cross sections exhibit uncertainties exceeding 10%, which must be addressed in order to accurately model positron fragmentation, particularly in the case of complex fragmentation reactions such as the production of ${}^{10}$ [34, 35]. This also impacts other Monte Carlo simulation platforms (such as FLUKA, MCNP and PHITS) which rely on accurate cross section data.
@@ 4 Conclusion
In this study, the accuracy with which Geant4 is able to predict the distribution of total positron annihilation yield and the distributions of individual positron-emitting fragmentation products (${}^{11}$, ${}^{10}$ and ${}^{15}$) during carbon or oxygen ion therapy was compared to experimental data. Three different hadronic inelastic physics models - Binary Ion Cascade (BIC), Quantum Molecular Dynamics (QMD) and Liege Intranuclear Cascade model (INCL) were used with ten different versions of Geant4 - 10.0.p04, 10.1.p03, 10.2.p03, 10.3.p03, 10.4.p03, 10.5.p01, 10.6.p03, 10.7.p02, 11.0 and 11.1, in three different homogeneous phantoms. The simulated and experimental data were compared using two different metrics - normalised mean squared error and the Pearson cross-correlation coefficient. Additionally, the differences between the simulated and experimental depth of maximum positron annihilation yield, as well as the distal point at which positron yield declines to 50% of the peak were evaluated. It was found that the accuracy of the hadronic inelastic physics models strongly depends on the version of Geant4 in which it was implemented, and newer versions of Geant4 were not always more accurate at predicting positron-emitting fragmentation compared to older versions. Furthermore, it was found that not all version/model combinations were able to satisfactorily predict the shape of positron annihilation or positron-emitting fragment distributions, even though they could provide a good estimation of the total positron annihilation yield and range. For future simulation studies of therapeutic irradiation using carbon or oxygen ion beams, it is recommended that Geant4 version 10.2 with the BIC model be used as it is currently the version/model combination best able to replicate the experimentally-observed total positron yield and the fragmentation product distributions, while the depth of the maximum positron yield and distal 50%-of-peak point were best predicted using the QMD model from Geant4 10.4, 10.6 (peak) and 10.2 (distal 50%-of-peak).

### v2

@@ 3.4 Overall recommendation
In the entrance region, BIC was clearly the best-performing model, with the best choice of Geant4 version depending on the particular metric and fragmentation product of interest. NMSE results generally favoured 10-10.4/BIC (especially 10.2/BIC), except for 10C yield, which was better predicted by 10.3+/INCL. Pearson CCC performance did not strongly favour any particular version/model combination, with at most 1/3 of test cases achieving the target CCC threshold of 0.95 for any version/model.
In the build-up and Bragg peak region and tail region, the results are more clear-cut. Results for the NMSE metric conclusively show that version 10.2/BIC is clearly the best choice for total positron yield, 11C production and 15O yield, while 10.5-11.1/INCL performed the best for 10C. Pearson CCC results are more mixed, but again, 10.2/BIC gives the best results for total positron annihilation yield, with most versions of Geant4 with BIC performing well. 10.6/QMD performed the best for 11C, 10/INCL and 10.1/INCL performed the best for 10C, and there was no clear winner for 15O.
Using the depth-of-maximum-yield metric, the smallest mean differences were obtained with 10.4/QMD and 10.6/QMD. These versions/models also achieved the equal-smallest maximum difference (-3 mm and +3 mm, respectively). Across all versions of Geant4, QMD demonstrated the best overall accuracy (lowest average mean difference in peak depth) and highest precision (lowest average standard deviation). INCL was the worst-performing model across all versions, with much larger maximum differences, and a consistent underestimation of depth of maximum yield across Geant4 versions, with the exception of version 11.1 (which, despite a mean difference of 0, exhibited a large standard deviation and maximum value). Standard deviations obtained using INCL were generally around double those of QMD and BIC. BIC also showed a consistent underestimation in depth of maximum yield, although the maximum differences were much smaller than for INCL. For context, the difference between the depth of the positron annihilation peak and the Bragg peak with monoenergetic ion beams is of the order of -5.6$\pm$0.8 mm for 12C and -6.6$\pm$0.8 mm for 16O [32, 33, 6].
@@
Finally, in the tail region, Geant4 10.2 with BIC and QMD again provided the best prediction of total positron and 11C yield in terms of NMSE, while 10.6/INCL performed the best for 10C and 15O. All version/model combinations performed well for total positron annihilation, 11C and 15O yield according to the Pearson CCC metric, while no version/model performed especially well for 10C.
@@
In the evaluation of individual positron-emitting fragment yield profiles, predictions of 10C distribution were generally the least accurate in terms of both the NMSE and Pearson CCC. Interestingly, the INCL model often performed the best for prediction of 10C fragment yield, although it rarely performed the best for total positron annihilation and 11C or 15O. Therefore, INCL should be considered for studies focusing on 10C fragmentation, with the caveat that range estimation will be less accurate with this model.
@@
Finally, it is worth noting that current evaluations of fragmentation cross sections exhibit uncertainties exceeding 10%, which must be addressed in order to accurately model positron fragmentation, particularly in the case of complex fragmentation reactions such as the production of 10C [34, 35]. This also impacts other Monte Carlo simulation platforms (such as FLUKA, MCNP and PHITS) which rely on accurate cross section data.
@@ 4 Conclusion
In this study, the accuracy with which Geant4 is able to predict the distribution of total positron annihilation yield and the distributions of individual positron-emitting fragmentation products (11C, 10C and 15O) during carbon or oxygen ion therapy was compared to experimental data. Three different hadronic inelastic physics models - Binary Ion Cascade (BIC), Quantum Molecular Dynamics (QMD) and Liege Intranuclear Cascade model (INCL) were used with ten different versions of Geant4 - 10.0.p04, 10.1.p03, 10.2.p03, 10.3.p03, 10.4.p03, 10.5.p01, 10.6.p03, 10.7.p02, 11.0 and 11.1, in three different homogeneous phantoms. The simulated and experimental data were compared using two different metrics - normalised mean squared error and the Pearson cross-correlation coefficient. Additionally, the differences between the simulated and experimental depth of maximum positron annihilation yield, as well as the distal point at which positron yield declines to 50% of the peak were evaluated. It was found that the accuracy of the hadronic inelastic physics models strongly depends on the version of Geant4 in which it was implemented, and newer versions of Geant4 were not always more accurate at predicting positron-emitting fragmentation compared to older versions. Furthermore, it was found that not all version/model combinations were able to satisfactorily predict the shape of positron annihilation or positron-emitting fragment distributions, even though they could provide a good estimation of the total positron annihilation yield and range. For future simulation studies of therapeutic irradiation using carbon or oxygen ion beams, it is recommended that Geant4 version 10.2 with the BIC model be used as it is currently the version/model combination best able to replicate the experimentally-observed total positron yield and the fragmentation product distributions, while the depth of the maximum positron yield and distal 50%-of-peak point were best predicted using the QMD model from Geant4 10.4, 10.6 (peak) and 10.2 (distal 50%-of-peak).